The difference between two numbers is 106. 6 times the smaller number is equal to 4 more than the larger. What are the numbers?

Answer by spashkin(2)

(X-Y = 106)

(6Y = X + 4)
-4 -4
X = 6Y-4
Move into other formula
6Y-4 -Y = 106
5Y -4 = 106
5Y = 110
Y= 22
Take this value to get X
X-22=106
X= 128
Bigger # = 128
Smaller # = 22